The product of the numbers on two consecutive pages in a book is 600. Find the page numbers.
x(x+1)=600
2x=600
x=300?

Answer by oberobic(2304)

Two consecutive page numbers would be, for example, 1 & 2 or 323 and 324 or x and x+1.
.
x(x+1) = 600
x^2 + x - 600 = 0
(x-24)(x+25) = 0
.
So there are two solutions: x=24 and x=-25.
.
But a negative page number does not make any sense. So the answer is:
.
x = 24
x+1= 25
.
Done